What is the concentration of the original solution if 450 mL, a certain solution is diluted in 750 mL water to form 2.0 L of 1.50 M HCl?

1
Expert's answer
2022-03-13T15:21:32-0400

n(HCl) = C(HCl) * V(solution) = 1.50 * 2.0 = 3.0 moles;

C0(HCl) = n(HCl)/V0(solution) = 3/0.45 = 6,67 М.

Answer: 6.67 M.